OpenStudy (agar.io):

What is the ninth term of the arithmetic sequence defined by the rule below? A(n)= -14+(n-1)(2) @TheSmartOne I got A(9) =2

TheSmartOne (thesmartone):

A(9) = -14 + (9-1)(2) A(9) = -14 + (8)(2) A(9) = -14 + 16 A(9) = 16 - 14 A(9) = 2 @Agar.io @~ You are correct! :)

OpenStudy (agar.io):

thnxs